rub, ย Home. my Tri-Cip· g: Mal-Labb, Mna..- (9 unread-dr-mro n Facebook m lyota FrancialSen ligher One Recent Unproctored Placement Assessment Due: Thursday 9:39 PM I Question 1 The perimeter of the rectangle below is 174 units. Find the length of side RS. Write your answer without variables. x +3 I Dont Know Submit ere to search DOLL
Answer #1

perimeter = 174 units

perimeter is sum of all sides

so, we can write the equation as

( 3x + 3 ) + ( 4x ) + ( 3x + 3) + ( 4x ) = 174

14x + 6 = 174

subtrcating 6 from both sides

14x = 174 - 6

14x = 168

dividing both sides by 14

x = 12

RS = 12 * 4

RS = 48

RS = 48